a mobile phone was listed for $432 less 30% 12.5% 5% . What is the exact single rate of discount that was allowed


1
Expert's answer
2021-10-10T16:47:38-0400

First "30" % discount:

"\frac{30}{100}\times432=129.6"

"432-129.6=302.4"

Second 12.5 % discount:

"\frac{12.5}{100}\times302.4=37.8"

"302.4-37.8=264.6"

Third 5% discount:

"\frac{5}{100}\times264.6=13.23"

"264.6-13.23=251.37"

The exact single rate of discount is:

"\frac{432-251.37}{432}\times100=41.81"

"=41.81" %